# mycapequ[][List of equations]

### v1

@@ V-A Methodology
The simulation methodology focuses in the optimization problem underlying our protocol definition. We tune the integer parameter $L$ in order to minimize the expectation of the coherence witness $W_{\mathrm{Fock}}$ over the state
$\theta_{N,\eta,\gamma,\lambda}=\Gamma_{\lambda}\Bigg{(}\Lambda_{\gamma}\bigg{(}\Xi_{\eta}\Big{(}\frac{1}{2}\sum_{i,j\in\{0,1\}}\mid i\rangle_{N}{}_{N}\langle j\mid\Big{)}\bigg{)}\Bigg{)},$ (20)
i.e., that one describing the selection of incoming pulses that Bob measures for computing such expectation. This leads to the definition of our cost function as [^1] Notice that, in practice, small dark-count effects allow to truncate the infinite sum appearing in Eq. (21) into just a few terms.
$\chi_{N,\eta,\gamma,\lambda}(l)\equiv\langle W_{\mathrm{Fock}}(l)\rangle_{\theta_{N,\eta,\gamma,\lambda}}=$ (21)
@@
$\vartheta_{\eta,\lambda,\gamma}(x,y,z)\equiv$ (22)
@@
Notice that minimizing $\chi_{N,\eta,\gamma,\lambda}(l)$ over $l$ yields the value of $L$ used at the implementation of the protocol. We propose to perform such minimization via brute-fore search, given the low dimension of the domain space.
@@ V-B Results
Fig. 2 shows, for three different noise levels also studied in Fig. 1, the dependency on $N$ for both the QBER obtained when reading out photon-number encoded bits and the optimal value of $\chi$ which, given any value below 0 and with a precision that increases with the amount of sent pulses, allows for detecting both eavesdropping and Alice-impersonation attacks. Remarkably, for both low-noise conditions and our estimated noise impact for a distance between parties of around $8$ km, the QBER rapidly decreases with $N$ while $\chi$ increases at a slower pace. Thus, for the indicated noise conditions, and for certain moderate values of $N$, our proposal enlarges the raw key and lowers its QBER with respect to single-photon-based implementation, while remaining secure.
Figure 2: For different levels of noise, optimal $\chi$ (left vertical axis) and expected QBER obtained with our employed bosonic qubit encoding (right vertical axis) as a function of $N$. Low, estimated and high noise characterizations coincide with those shown in Fig. 1.
@@ VI Conclusions and further research
Quantum Key Growing (QKG) is of critical importance in the current landscape of quantum communications, both due to its theoretical and experimental progress and its essential role in countering emerging quantum threats [54, 55]. In this work, we have proposed a novel Hybrid Key Growing (HKG) protocol based on bipartite encoding, coherence witnessing, and a classical assumption on the physical layer. As shown in Sections IV and V-B, the protocol is compatible with real-world implementation under moderate noise conditions.
Although mature solutions such as decoy-state QKG remain highly effective and widely adopted, our approach explores alternative strategies with distinct operational principles. We believe that such diversity is vital to advancing in the field. Future research will examine whether the adversarial model used in our analysis poses challenges to existing protocols and whether our method improves scalability in networked settings, e.g., in Hybrid Conference Key Growing (HCKG).

### v2

@@ V-A Methodology
The simulation methodology focuses in the optimization problem underlying our protocol definition. We tune the integer parameter $L$ in order to minimize the expectation of the coherence witness $\hat{W}_{\mathrm{Fock}}$ over the state
$),\theta_{N,\eta,\gamma,\lambda}=\Gamma_{\lambda}\Bigg(\Lambda_{\gamma}\bigg(\Xi_{\eta}\Big(\frac{1}{2}\sum_{i,j\in\{0,1\}}\mid i\rangle_{N}{}_{N}\langle j\mid\Big)\bigg$$\Bigg)$, (20)
i.e., that one describing the selection of incoming pulses that Bob measures for computing such an expectation. This leads to the definition of our cost function as [^1] Notice that, in practice, small dark-count effects allow to truncate the infinite sum appearing in Eq. (21) into just a few terms.
$\chi_{N,\eta,\gamma,\lambda}(l):=\langle\hat{W}_{\mathrm{Fock}}(l)\rangle_{\theta_{N,\eta,\gamma,\lambda}}=$ (21)
@@
$\vartheta_{\eta,\lambda,\gamma}(x,y,z):=$ (22)
@@
Notice that minimizing $\chi_{N,\eta,\gamma,\lambda}(l)$ over $l$ yields the value of $L$ used at the implementation of the protocol. We propose to perform such a minimization via brute-fore search, given the low dimension of the domain space.
@@ V-B Results
Fig. 2 shows, for three different noise levels also studied in Fig. 1, the dependency on $N$ for both the QBER obtained when reading out photon-number-encoded bits and the optimal value of $\chi$. Remarkably, for both low-noise conditions and our estimated noise impact for a distance between parties of around $8$ km, the QBER rapidly decreases with $N$ while $\chi$ increases at a slower pace. Thus, for the indicated noise conditions, and for certain moderate values of $N$, our proposal enlarges the generated raw key and lowers its QBER with respect to single-photon-based implementation, while remaining secure.
Figure 2: For various noise levels, the optimal $\chi$ (left axis) and expected QBER with our bosonic qubit encoding (right axis) are shown as a function of $N$. Noise characterizations match those in Fig. 1.
@@ VI Conclusions and further research
Quantum Key Growing (QKG) is of critical importance in the current landscape of quantum communications, both due to its theoretical and experimental progress and its essential role in countering emerging quantum threats [53, 54]. In this work, we have proposed a novel Hybrid Key Growing (HKG) protocol based on bipartite encoding, coherence witnessing, and a classical assumption on the physical layer. As shown in Sections IV and V-B, the protocol is compatible with real-world implementation under moderate noise conditions, assuming that our encoding-and-decoding scheme is plausible.
Although mature solutions such as decoy-state QKG remain highly effective and widely adopted, our approach explores alternative strategies with distinct operational principles. We believe that such a diversity is vital to advancing in the field. Future research will examine whether the adversarial model used in our analysis poses challenges to existing protocols and whether our method improves scalability in networked settings, e.g., in Hybrid Conference Key Growing.
Finally, as we became aware during the course of our research, we notice that generating and detecting states in the Fock basis remains a significant challenge with current technology. The recent work presented in [55] offers a promising direction for overcoming these difficulties.